# MTQA: Matrix of Thought for Enhanced Reasoning in Complex Question Answering

### v1

@@ Abstract
Complex Question Answering (QA) is a fundamental and challenging task in NLP. While large language models (LLMs) exhibit impressive performance in QA, they suffer from significant performance degradation when facing complex and abstract QA tasks due to insufficient reasoning capabilities. Works such as Chain-of-Thought (CoT) and Tree-of-Thought (ToT) aim to enhance LLMs’ reasoning abilities, but they face issues such as in-layer redundancy in tree structures and single paths in chain structures. Although some studies utilize Retrieval-Augmented Generation (RAG) methods to assist LLMs in reasoning, the challenge of effectively utilizing large amounts of information involving multiple entities and hops remains critical. To address this, we propose the Matrix of Thought (MoT), a novel and efficient LLM thought structure. MoT explores the problem in both horizontal and vertical dimensions through the ”column-cell communication” mechanism, enabling LLMs to actively engage in multi-strategy and deep-level thinking, reducing redundancy within the column cells and enhancing reasoning capabilities. Furthermore, we develop a fact-correction mechanism by constructing knowledge units from retrieved knowledge graph triples and raw text to enhance the initial knowledge for LLM reasoning and correct erroneous answers. This leads to the development of an efficient and accurate QA framework (MTQA). Experimental results show that our framework outperforms state-of-the-art methods on four widely-used datasets in terms of F1 and EM scores, with reasoning time only 14.4% of the baseline methods, demonstrating both its efficiency and accuracy. The code for this framework is available at https://github.com/lyfiter/mtqa.
Introduction
The human transition from learning to practice generally entails comprehending and reasoning over diverse information sources, followed by making decisions that align with—or even extend beyond—the original information. Recent studies have demonstrated that large language models (LLMs) can replicate this learning-to-practice paradigm in complex domains such as intelligence analysis (Wu et al. 2024) and creative writing (Qin et al. 2024). Complex question answering (QA) is a foundational and critical downstream task for LLMs (Zhuang et al. 2023); it demands that a QA system command extensive knowledge and perform sophisticated reasoning over that knowledge to deliver satisfactory answers.
@@ Introduction
Figure 1: Comparison of Chain of Thought, Thought of Tree, and MoT for Complex Question Answering.
Recently, large language models (LLMs) have become one of the most promising and efficient solutions in the QA field, thanks to their massive pre-trained knowledge and the reasoning abilities they autonomously generate during training(Guo et al. 2025). However, some studies indicate that LLMs exhibit significant shortcomings in complex question-answering reasoning tasks. Various research efforts have attempted to enhance LLMs’ reasoning performance using thought structures. For example, the Chain-of-Thought (CoT) framework(Wei et al. 2022) breaks down the reasoning process of LLMs into a series of ordered steps, thereby improving their reasoning capabilities. However, CoT suffers from a single reasoning path, limiting its ability to explore multiple reasoning paths within the solution space, and may also result in incorrect ideas that cannot be corrected in time, leading to deviations. Tree-of-Thought (ToT)(Yao et al. 2023a) organizes the reasoning process into a hierarchical tree structure, allowing the model to reason across multiple layers and branches. However, its node selection mechanism is inefficient and often produces redundant nodes, leading to resource wastage and ineffective reasoning.
@@
Furthermore, some studies have employed Retrieval-Augmented Generation (RAG) to assist LLMs in answering questions, enhancing them with external knowledge through retrieval and generation, leading to more accurate and reliable answers(Cuconasu et al. 2024; Fan et al. 2024). Currently, mainstream reasoning methods incorporate RAG into thought structures to provide LLMs with the ability to validate facts. RAT(Wang et al. 2024) refines the step-by-step reasoning process of CoT by adding a RAG correction at each reasoning stage to ensure the reasoning trajectory remains on track. RATT(Zhang et al. 2025) combines the multi-branch, multi-level reasoning of tree structures with RAG’s external knowledge to correct factual errors, thereby strengthening LLMs’ reasoning capabilities. However, these works use vanilla RAG as the fact-validation module, which limits its ability to assist in reasoning and correct facts. In complex question-answering tasks, especially in open-ended tasks involving multiple entities and multi-hop information, the retrieved knowledge often contains irrelevant or even erroneous information, which misguides the LLMs’ reasoning process. The quality of the factual verification knowledge provided by RAG becomes one of the key factors in guiding LLMs’ reasoning.
To address the aforementioned issues and improve the reliability and accuracy of LLMs in complex question answering tasks, we aim to develop a method that efficiently utilizes existing knowledge and performs multi-branch reasoning with low redundancy to generate accurate answers. In terms of efficiently utilizing existing knowledge, we focus on exploring the representation of knowledge retrieved by RAG to assist LLMs in answering questions. In the context of multi-branch and low-redundancy reasoning, we aim for the model to autonomously adopt different perspectives to understand the existing knowledge and perform reasoning that leads to an answer close to the correct one, all while using the same initial knowledge.
@@
To achieve these goals, we propose a novel Matrix of Thought (MoT) approach to ensure that the reasoning process for complex question-answering tasks is multi-branching and low-redundancy. As shown in Figure 1, in MoT, our RAG retrieves entity-relation-entity triples that match the question, along with their corresponding original text segments, which are then constructed into knowledge units to serve as auxiliary knowledge for initiating reasoning and fact verification. Additionally, we introduce a column cell communication mechanism that actively helps the LLM explore alternative perspectives for solving the problem, thereby reducing the redundancy in the generated reasoning process within the column. Then, MoT combines the various branching strategies with RAG’s auxiliary knowledge for summarization and optimization, using this as the reasoning initiation knowledge for the next column cell, thereby approaching the most promising answer.
@@
1. We propose an innovative large model reasoning paradigm, MoT, which addresses the issues of information redundancy within the same layer of tree structures and the single decision path in chain structures. MoT enhances the flexibility and adaptability of reasoning at a minimal cost.
2. We have developed a novel framework for complex QA tasks, the Matrix of Thought Question Answering (MTQA) framework. This framework innovatively combines KG triples with their corresponding original text into knowledge units for knowledge enhancement and correction. By integrating the MoT reasoning paradigm, it achieves a fully automated QA pipeline, enabling efficient reasoning on the problem from both breadth and depth.
@@
3. We conducted a series of experiments to validate the efficiency and reliability of our method in complex question answering across various domains. The results demonstrate that our method holds a clear advantage over existing approaches.
Related Work
Thought Structures for LLMs
@@
The thought structures proposed (Ding et al. 2023; Minaee et al. 2024) is a mainstream prompt engineering approach that guides LLMs to engage in more logical thought processes, thereby improving the quality of generated responses. Chain of Thought (Wei et al. 2022) is a representative work that guides LLMs to generate intermediate reasoning steps within the prompt, thereby steering the large model through the reasoning process. Wang et al. developed the Self-consistency with Chain of Thought (CoT-SC) method (Wang et al. 2023), which extends CoT by solving a problem using multiple independent and distinct chains of thought. The most reliable answer is then selected as the final output response. The Tree of Thoughts (ToT) (Yao et al. 2023a) abstracts the reasoning process of a problem into a tree search. By constructing multiple branches and evaluating their contributions to solving the problem, ToT guides the LLM in reasoning, enabling a broad and global exploration of the solution space.
@@
Current mainstream research focuses on CoT and ToT; however, the inherent properties of their structures present performance bottlenecks, leading to issues such as single reasoning paths, generation of redundant information, and the risk of deviating from the reasoning trajectory due to incorrect ideas during inference. One of the key focuses of our research is to develop a more efficient and reliable thought structure paradigm, enabling LLMs to perform exceptionally well in QA tasks and beyond within this new framework.
Retrieval-Augmented Generation (RAG) for LLM on QA
@@
When LLMs solve extremely complex tasks, they suffer from ”complete accuracy collapse”(Shojaee*† et al. 2025), significant performance degradation, and LLMs hallucination problems due to insufficient reasoning capabilities(Sriramanan et al. 2024). Retrieval-Augmented Generation (RAG)(Lewis et al. 2020) is a pivotal technique for mitigating such hallucinations and enhancing the quality of LLM outputs. Retrieval-Augmented Generation (RAG) stores task-relevant knowledge in an external repository and, when the LLM is invoked, retrieves pertinent information to serve as supplementary context. GraphRAG (Edge et al. 2024) extends this paradigm by incorporating knowledge graphs (KGs): an LLM produces concise community-level summaries that both abstract the input documents and provide graph indices; each community then generates its response independently, and the individual outputs are subsequently aggregated into a global answer. Recent studies have begun to integrate explicit reasoning paradigms into the RAG workflow. Retrieval-Augmented Thought (RAT)(Wang et al. 2024) merges the RAG retrieval step directly into the chain-of-thought process, allowing the model to iteratively refine and correct its reasoning step by step. Retrieval-Augmented Thought Tree (RATT) (Zhang et al. 2025) adopts a tree-structured reasoning framework in which each node invokes RAG for factual verification, thereby enhancing the logical soundness and factual reliability of LLM reasoning.
@@
Methodology
This section presents the Matrix of thought Question Answering (MTQA) framework in detail. As illustrated in Figure 2, MTQA is constructed atop the novel MoT reasoning paradigm to tackle complex QA tasks. The framework is designed to make fuller use of existing knowledge and to equip LLMs with both breadth and depth when exploring the problem’s solution space. Our method improves the reliability and comprehensiveness of LLM-generated answers from two aspects: adopting a superior knowledge representation to enhance the LLM’s comprehension. And employing a powerful thought structures to strengthen the LLM’s reasoning capability. The framework consists of two parts: (1) retrieval-augmentation, (2) Matrix of thought.
Figure 2: Framework of MTQA. Given an input including a query and source documents, the LLM retrieves relevant knowledge units from both knowledge graphs and text databases. These knowledge units are then integrated into the Matrix of Thought (MoT), where multiple reasoning strategies are explored in parallel through column-cell communication. The reasoning process is enhanced by retrieval-augmented knowledge, and the final answer is refined through correction and summarization.
Problem Definition
@@
In MTQA, our objective is to construct a matrix of thought $M$ that can carry out the widest and deepest possible inference within the solution space of question $Q$, thereby identifying the optimal answer $A$.
Retrieval-Augmented
Step 1: Knowledge Database construction.In MTQA, the first step is to build a retrievable external knowledge base and partition the given set of documents into chunks whose size is suitable for processing by the LLM:
$C=\{c_{1},\,c_{2},\,c_{3},\,\dots,\,c_{k}\},\quad k=\left\lceil\frac{N_{D}}{C_{\text{len}}}\right\rceil$ (1)
@@
$L_{c_{i}}=\operatorname{embed}\!\bigl{(}\operatorname{template}(C_{i},x)\bigr{)}$ (2)
@@
$v,\;\varepsilon=\text{Extraction \& identification}\!\left(\sum_{i=1}^{k}L_{c_{i}}\right)$ (3)
Subsequently, we develop guidance llm to construct a key-value pair $(k,v)$ for each recognized and extracted entity $v$ and the relation $\varepsilon$ between them. The keys are composed of individual words or short phrases and are later used as query terms to retrieve the corresponding entities and relations. For entities, their names function as keys; for relations, the LLM summarizes the text that describes the link between two entities to generate multiple keys. The associated values are the source passages in the external knowledge base from which those entities and relations are drawn. The key–value mechanism both accelerates retrieval and preserves direct access to the source passages in the external knowledge base. During the generation of entities and their corresponding relations, the LLM augments each item with an explanatory description, allowing subsequent LLM-based reasoning modules to exploit this enriched knowledge more effectively for refinement and enhancement.
$\hat{v},\;\hat{\varepsilon}=\operatorname{Deduplication}\!\bigl{(}v,\;\varepsilon\bigr{)}$ (4)
After completing the extraction of entity $v$ and relation $\varepsilon$, we perform a de-redundancy operation to identify and merge the duplicates of $v$ and $\varepsilon$ to obtain the unique $\hat{v}$ and $\hat{\varepsilon}$ in the whole knowledge graph, which improves the overhead of the subsequent operations on the knowledge graph and ensures that there is no redundancy in the retrieved knowledge.
$G=\bigl{(}\hat{v},\;\hat{\varepsilon}\bigr{)}$ (5)
@@
The key-value structure of the knowledge graph that we construct as the external knowledge base for MTQA offers two advantages:
(1) Compared to indexing the original text passages, the knowledge graph stores structural information, providing the MTQA framework with multi-hop reasoning capabilities to integrate global information, enabling LLMs to understand knowledge across text blocks.
(2) The key-value structure strongly associates the knowledge graph with its source texts, facilitating efficient retrieval while enabling the construction of knowledge units that combine knowledge graph triples with source text-enhanced knowledge.
Step 2: Retrieval knowledge through queries. After constructing the external knowledge base, we build the retrieval function mechanism to match the knowledge that best supports the query. We use a dual retrieval mechanism (Edge et al. 2024) to extract relevant information from the numerous and complex intertwined dependencies of entities in the external knowledge base. This includes detailed queries and summary queries. Detailed queries focus on specific details and require precise identification of relevant entities or relations for factual judgment or providing key evidence. Summary queries, on the other hand, are broader in scope, summarizing topics or abstracts. They typically involve searching across multiple entity-relation triples or even text blocks, and are used to answer open-ended or comprehensive questions. We perform retrieval by combining graph structure information with vector representations. The retrieval process is mainly divided into three steps:
(1) Extract the keywords from the query $Q$, distinguishing between detailed keywords $k^{(d)}$ and generalized keywords $k^{(g)}$.
(2) Match the extracted detailed keywords with the entity keys in the vector database, and match the generalized keywords with the relation keys.
(3) Further retrieve the neighboring nodes (one-hop and two-hop adjacent nodes) in the knowledge graph for the matched entities and relations.
The retrieved knowledge consists of entities (nodes of the knowledge graph), relations (edges of the knowledge graph), and the original text (indexed by the key-value pairs of nodes and edges), forming knowledge units, as shown in Figure 2. The knowledge units simultaneously contain structured information, supporting multi-hop reasoning and emphasizing key information through knowledge graph triples(Saxena, Tripathi, and Talukdar 2020), as well as undistorted, broader content original (Lewis et al. 2020) (such as complex metaphors, emotional expressions, etc.), which better aligns with the preferences of LLMs (Dong et al. 2025) in original text. Experimental results demonstrate that these knowledge units significantly enhance the reasoning performance of LLMs.
Matrix of Thought

### v2

@@ Abstract.
Large Language Models (LLMs) face significant accuracy degradation due to insufficient reasoning ability when dealing with complex and abstract tasks. Thought structures such as Chain of Thought (CoT) and Tree of Thought (ToT) focus on enhancing the reasoning capability of LLMs. However, they suffer from inherent drawbacks such as redundancy within the same layer of the tree structure and the singularity of the paths in the chain structure. Some studies have utilized Retrieval-Augmented Generation (RAG) methods to enhance CoT and ToT in mitigating hallucinations in LLMs, yet the fundamental shortcomings of the thought structures still persist. Furthermore, when dealing with multi-entity and multi-hop information, the retrieved verification knowledge often contains large amounts of fragmented, superficial, or even erroneous data, misleading the reasoning process of LLMs. To address these issues, we propose the Matrix of Thought (MoT), a novel and efficient thought structure for LLMs. MoT explores problems in both horizontal and vertical dimensions through a ”column-cell communication” mechanism, enabling LLMs to actively engage in multi-strategy and deep thinking while reducing redundancy in the thought nodes within the column cells, thereby enhancing the reasoning capability of LLMs. Additionally, through a fact-correction mechanism, it leverages the knowledge graph triples retrieved by RAG and the original text to construct knowledge units and correct erroneous answers. To validate the effectiveness of this method, we conducted extensive experiments in three tasks: 24-point game, question answering evaluation, and proposition writing. The results demonstrate that our framework outperforms state-of-the-art methods, with reasoning time only 14.4% of that of the baseline method, proving its efficiency and accuracy.
LLM Reasoning, Thought Structure, RAG, Matrix of Thought (MoT)
isbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
The process by which humans move from learning to practice often involves understanding and reasoning through various pieces of information, leading to decisions that align with or even go beyond the content described in the original information. Recent studies have shown that Large Language Models (LLMs) can achieve a similar process of learning to practice in complex fields such as intelligence analysis(Wu et al., 2024) and creative writing(Qin et al., 2024). However, because they handle problems and generate thoughts through token-level predictions, this approach limits their ability to expand high-level or multi-perspective reasoning(Huang et al., 2022).
@@ 1. Introduction
Figure 1. Comparison of Chain of Thought, Thought of Tree, and MoT for Complex Question Answering.
Some studies attempt to enhance the reasoning performance of LLMs through thought structures. For example, Chain-of-Thought (CoT) (Wei et al., 2022), as shown in Figure 1, decomposes the reasoning process of LLMs into a series of ordered reasoning steps, thereby enhancing its reasoning capability. However, its reasoning path is singular, unable to explore multiple paths within the solution space, and errors in reasoning may not be corrected in time, leading to deviations. Tree-of-Thought (ToT) (Yao et al., 2023a) structures the reasoning process into a hierarchical tree model, allowing the model to reason across multiple levels and branches. However, its node selection mechanism is not efficient, and it generates a large number of redundant nodes, resulting in resource waste and inefficient reasoning.
@@
As the complexity of tasks increases, methods such as CoT and ToT are prone to deviations from the facts during the reasoning process due to hallucinations [citation], which cannot be corrected in time, leading to a misalignment in the subsequent reasoning process. Some works employ Retrieval-Augmented Generation (RAG) to guide and correct the responses of LLMs by retrieving external knowledge to supplement the models, followed by generation to produce more accurate and reliable answers(Cuconasu et al., 2024; Fan et al., 2024).Currently, mainstream reasoning methods embed RAG into thought structures to provide LLMs with the ability to verify facts. RAT(Wang et al., 2024) breaks down the stepwise reasoning process of CoT and adds the RAG correction process to each intermediate reasoning step to ensure that the reasoning trajectory does not deviate. RATT(Zhang et al., 2025) combines the multi-branch, multi-level reasoning of the tree structure with RAG’s external knowledge to correct factual errors, thereby enhancing the reasoning ability of LLMs. However, in the aforementioned methods, RAG simply supports knowledge based on similarity matching, and the retrieved verification knowledge often lacks coherence, creating fragmentation and misalignment with the LLM’s preferences. Additionally, when performing complex question-answering tasks, especially open-domain tasks involving multiple entities and multi-hop information, the retrieved verification knowledge often contains large amounts of irrelevant or even erroneous information, thus misleading the reasoning process of LLMs.
@@
To achieve the aforementioned goal, we propose a novel Matrix of Thought (MoT) method, which ensures that the reasoning process can proceed with multiple branches and low redundancy in the context of solving complex problems. As shown in Figure 1, in MoT, we use RAG to retrieve KG triples and their corresponding original text segments that match the question, which are then integrated as auxiliary knowledge for initiating reasoning and fact verification. Additionally, we introduce a column communication mechanism that actively helps the LLM explore other perspectives for solving the problem, thereby reducing redundancy in the generated reasoning process within the column. Subsequently, MoT combines the various branch strategies with the auxiliary knowledge from RAG, summarizes and optimizes them, and uses this as the reasoning initiation knowledge for the next column unit, thereby converging towards the most promising answer.
@@
  • We propose an innovative thought structure paradigm, MoT, for large model reasoning. It addresses the issues of information redundancy within the same layer of tree structures and the singularity of decision paths in chain structures during reasoning, enhancing the flexibility and adaptability of reasoning with minimal cost.
  • We explore a novel method of embedding RAG within the LLM thought structure, constructing knowledge units to replace the conventional naive knowledge in RAG, in order to align with the LLM’s understanding preferences and enhance the integrity of the retrieved knowledge.
@@
  • We conducted a series of experiments to validate the efficiency and reliability of our method in various complex tasks. The results demonstrate that our method holds a clear advantage over existing approaches.
2. Related Work
2.1. Thought Structures for LLMs
@@
The thought structures proposed (Ding et al., 2023; Minaee et al., 2024) is a mainstream prompt engineering approach that guides LLMs to engage in more logical thought processes, thereby improving the quality of generated responses. Chain of Thought (Wei et al., 2022) is a representative work that guides LLMs to generate intermediate reasoning steps within the prompt, thereby steering the large model through the reasoning process. Wang et al. developed the Self-consistency with Chain of Thought (CoT-SC) method (Wang et al., 2023), which extends CoT by solving a problem using multiple independent and distinct chains of thought. The most reliable answer is then selected as the final output response. The Tree of Thoughts (ToT) (Yao et al., 2023a) abstracts the reasoning process of a problem into a tree search. By constructing multiple branches and evaluating their contributions to solving the problem, ToT guides the LLM in reasoning, enabling a broad and global exploration of the solution space.
@@
Current mainstream research focuses on CoT and ToT; however, the inherent properties of their structures present performance bottlenecks, leading to issues such as single reasoning paths, generation of redundant information, and the risk of deviating from the reasoning trajectory due to incorrect ideas during inference. One of the focuses of our research is to develop a more efficient and reliable thought structure paradigm, enabling the LLM to achieve superior performance in various complex tasks within this paradigm.
2.2. Retrieval-Augmented Generation (RAG) for LLM
@@
When LLMs solve extremely complex tasks, they suffer from ”complete accuracy collapse”(Shojaee*† et al., 2025), significant performance degradation, and LLMs hallucination problems due to insufficient reasoning capabilities(Sriramanan et al., 2024). Retrieval-Augmented Generation (RAG)(Lewis et al., 2020) is a pivotal technique for mitigating such hallucinations and enhancing the quality of LLM outputs. Retrieval-Augmented Generation (RAG) stores task-relevant knowledge in an external repository and, when the LLM is invoked, retrieves pertinent information to serve as supplementary context. GraphRAG (Edge et al., 2024) extends this paradigm by incorporating knowledge graphs (KGs): an LLM produces concise community-level summaries that both abstract the input documents and provide graph indices; each community then generates its response independently, and the individual outputs are subsequently aggregated into a global answer. Recent studies have begun to integrate explicit reasoning paradigms into the RAG workflow. Retrieval-Augmented Thought (RAT)(Wang et al., 2024) merges the RAG retrieval step directly into the chain-of-thought process, allowing the model to iteratively refine and correct its reasoning step by step. Retrieval-Augmented Thought Tree (RATT) (Zhang et al., 2025) adopts a tree-structured reasoning framework in which each node invokes RAG for factual verification, thereby enhancing the logical soundness and factual reliability of LLM reasoning.
@@
3. Methodology
This section presents the MoT framework in detail. As illustrated in Figure 2. The framework is designed to make fuller use of existing knowledge and to equip LLMs with both breadth and depth when exploring the problem’s solution space. The framework consists of two parts: (1) retrieval-augmentation, (2) Matrix of thought.
Figure 2. Framework of MoT. Given an input including a query and source documents, the LLM retrieves relevant knowledge units from both knowledge graphs and text databases. These knowledge units are then integrated into the Matrix of Thought (MoT), where multiple reasoning strategies are explored in parallel through column-cell communication. The reasoning process is enhanced by retrieval-augmented knowledge, and the final answer is refined through correction and summarization.
3.1. Problem Definition
@@
In MoT, our objective is to construct a matrix of thought $M$ that can carry out the widest and deepest possible inference within the solution space of question $Q$, thereby identifying the optimal answer $A$.
3.2. Retrieval-Augmented
Step 1: Knowledge Database construction.In MoT, the first step is to build a retrievable external knowledge base and partition the given set of documents into chunks whose size is suitable for processing by the LLM:
(1) $C=\{c_{1},\,c_{2},\,c_{3},\,\dots,\,c_{k}\},\quad k=\left\lceil\frac{N_{D}}{C_{\text{len}}}\right\rceil$
@@ 3.2. Retrieval-Augmented
(2) $L_{c_{i}}=\operatorname{embed}\!\bigl(\operatorname{template}(C_{i},x)\bigr)$
@@
(3) $v,\;\varepsilon=\text{Extraction \& identification}\!\left(\sum_{i=1}^{k}L_{c_{i}}\right)$
Subsequently, we develop guidance llm to construct a key-value pair $(k,v)$ for each recognized and extracted entity $v$ and the relation $\varepsilon$ between them(Guo et al., 2024).
(4) $\hat{v},\;\hat{\varepsilon}=\operatorname{Deduplication}\!\bigl(v,\;\varepsilon\bigr)$
After completing the extraction of entity $v$ and relation $\varepsilon$, we perform a de-redundancy operation to identify and merge the duplicates of $v$ and $\varepsilon$ to obtain the unique $\hat{v}$ and $\hat{\varepsilon}$ in the whole knowledge graph.
[…]
@@
Step 2: Retrieval knowledge through queries. After constructing the external knowledge base, we build the retrieval function mechanism to match the knowledge that best supports the query. We use a dual retrieval mechanism (Edge et al., 2024) to extract relevant information from the numerous and complex intertwined dependencies of entities in the external knowledge base.
The retrieved knowledge consists of entities (nodes of the knowledge graph), relations (edges of the knowledge graph), and the original text (indexed by the key-value pairs of nodes and edges), forming knowledge units, as shown in Figure 2. The knowledge units simultaneously contain structured information, supporting multi-hop reasoning and emphasizing key information through knowledge graph triples(Saxena et al., 2020), as well as undistorted, broader content original (Lewis et al., 2020) (such as complex metaphors, emotional expressions, etc.), which better aligns with the preferences of LLMs (Dong et al., 2025) in original text. Experimental results demonstrate that these knowledge units significantly enhance the reasoning performance of LLMs.
3.3. Matrix of Thought